OpenStudy (anonymous):

Find the 11th term of the arithmetic sequence. a_3 = 20.5 and a_8 = 13, a_11 = ______.

OpenStudy (anonymous):

Could you show us wat you've done so far with the question please?

OpenStudy (anonymous):

i dont know how to do it

jimthompson5910 (jim_thompson5910):

a_8 = (a_3) + 5d 13 = 20.5 + 5d solve for d

OpenStudy (anonymous):

Okay, well, let's start wiht the formula of an arithmetic sequence. what's the formula?

OpenStudy (anonymous):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

Okay, let's have this arithmetic sequence: \[a_{1}+(a_{1}+d)+(a_{1}+2d)+(a_{1}+3d)...\]

OpenStudy (anonymous):

AS you keep going up to the 8th term, what would the expression be for the 8th term?

OpenStudy (anonymous):

So basically what would: \[a_{8}=?\] Using your formula for an arithmetic sequence in terms of a and d?

OpenStudy (anonymous):

@tomtom777 Are you there?

OpenStudy (anonymous):

hello @Azteck sorry i had dinner and idk understand

OpenStudy (anonymous):

What do you mean? You just substitute n with 8 and get an expression in terms of a and d.

OpenStudy (anonymous):

what about a1

OpenStudy (anonymous):

Does not matter. it will cancel out afterwards.

OpenStudy (anonymous):

a1 or a. it's the same thing.

OpenStudy (anonymous):

The formula can be written as this: \[T_{n}=a+(n-1)d\]

OpenStudy (anonymous):

I'm used to it being written as that, so you might have to adapt or just use common sense that when I say a it means a_{1} etc. ;)

OpenStudy (anonymous):

can u just solve it

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):

I'm not going to gain anything, you're not going to gain anything. Wouldn't it be better to not ask about the question at all if you want us to do the question for you? We're here to HELP and GUIDE you so that you know how to do it and everybody will be happy.

OpenStudy (anonymous):

So you know the formula for an arithmetic term. \[T_{n}=a+(n-1)d\] What would T_{8} be? \[T_{8}=a+(8-1)d\]